OpenStudy (anonymous):

What is the slope of the line through the points (2,−5) and (3,−8)?

OpenStudy (anonymous):

i dont get it whats the answer so i can find the right way to do it

OpenStudy (anonymous):

slope formula is\[\frac{ y2-y1 }{ x2-x1}\]

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

Forgot to mention that lool

OpenStudy (anonymous):

y2-y1 / x2- x1 = m

OpenStudy (anonymous):

Yeah read it wrong, all you need is that formula.

OpenStudy (anonymous):

but whats the answer ? cuz it must be something like that 2/3 or 2-/3 cuz i am doing homework on a website

OpenStudy (anonymous):

-8 - (-5) / 3- 2 = ?

OpenStudy (anonymous):

-3 / 1 = -3

OpenStudy (anonymous):

so its -3?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

love u :D

OpenStudy (anonymous):

:)